OpenStudy (ttop0816):

cosθ(tanθ + cotθ) = _____ 1 cos^2θ << correct?? cscθ secθ

OpenStudy (anonymous):

1/sin(θ) = csc (θ)

OpenStudy (anonymous):

thats the answer :)

OpenStudy (solomonzelman):

\[\color{blue}{ \cosθ(\tanθ + \cotθ) }\] \[\color{blue}{ \cosθ( \frac{\sinθ}{\cosθ} + \frac{\cosθ}{\sinθ}) }\]expand,\[\color{blue}{ \sinθ+ \frac{\cos^2θ}{\sinθ} }\] \[\color{blue}{ \sinθ+ \frac{1-\sin^2θ}{\sinθ} }\] \[\color{blue}{ \sinθ+ \frac{1}{\sinθ} - \frac{\sin^2θ}{\sinθ} }\] \[\color{blue}{ \sinθ+ \cscθ -\sinθ }\] \[\color{blue}{ \cscθ }\] clear?

OpenStudy (anonymous):

hey why this Equation thing not working ?

OpenStudy (solomonzelman):

Idk, you are having some problems with the editor. Well, for me it's working well.

OpenStudy (anonymous):

dont know man, I can't even see the equation

OpenStudy (solomonzelman):

I know what you are seeing, \color{blue}{ sinθ+ cscθ -sinθ } that?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

Yeah, I initially thought it was because of the connection the browser and all that. On my computer latex used to slow don't and show as symbols like ones I showed you, not like equations, but I just fixed my connection to work instantly and latex worked. I don't think that you are having that problem though, dk what could it exactly be, not good with computers.

OpenStudy (anonymous):

hmm I'll ask some system admin might be a bug in the system cause latex works well with other sites :) appreciate your help

OpenStudy (solomonzelman):

Yeah, ask a moderator.

OpenStudy (anonymous):

will do